A train starts from rest at a station and travels with a constant acceleration of 1m/s2. Determine the velocity of the train when t = 30s and the distance travelled during this time.
1
Expert's answer
2020-08-25T11:14:33-0400

The velocity of a train

"v=v_0+at\\
=0\:\rm m/s+1\: m/s^2\times 30\: s=30\: m/s"

The distance travelled by a train

"d=v_0t+\frac{at^2}{2}\\
=0\:\rm m/s\times 30\: m/s+\frac{1\: m/s^2\times (30\: s)^2}{2}=450\: m"